In an investigation of toxins produced by molds that infect corn crops, a biochemist prepares extracts of the mold culture with organic solvents and then measures the amount of the toxic substance per gram of solution. From 13 preparations of the mold culture, the following 90% confidence interval for the average amount of the toxic substance per gram of solution was produced, (1.31, 1.6). What was the sample standard deviation? (2 decimals)
Answer #1

Solution :

Sample mean = = (1.31 + 1.6) / 2 = 1.455

Margin of error = E = Upper confidence interval -   = 0.145

t/2,df = 1.782

s = E * n / t/2,df

= 0.145 * 13 / 1.782

= 0.29

Sample standard deviation = 0.29